# Nonlinear Dynamics of Kink Configurations: From Small to Large Kink Collisions

### v1

@@ III.2 Large Kinks Scattering
Fig. 6 illustrates the scattering of large kinks for different values of $\lambda$ and initial velocities, with initial positions $X_{0k}=-X_{0ak}=-10$ in all cases. In Fig. 6, for $\lambda=0.2$ and $v_{0k}=-v_{0ak}=0.1$, the collision results in the creation of two small kinks, which move away symmetrically after the interaction. Increasing $\lambda$ to 0.6 with the same initial velocity ($v_{0k}=-v_{0ak}=0.1$), as shown in Fig. 6, also leads to the creation of two small kinks, but with higher velocities, indicating that a larger $\lambda$ enhances the energy transfer to the resulting small kinks. For $\lambda=0.8$ and $v_{0k}=-v_{0ak}=0.1$ (Fig. 6), the collision produces two pairs of small kinks, suggesting that higher $\lambda$ values facilitate the creation of additional small kinks due to changes in the potential $U_{\lambda}(\chi)$. A similar outcome is observed in Fig. 6 for $\lambda=0.6$ but with a higher initial velocity ($v_{0k}=-v_{0ak}=0.9$), where two pairs of small kinks are created, highlighting the role of initial velocity in increasing the number of small kinks produced. For $\lambda=0.9$ and $v_{0k}=-v_{0ak}=0.2$ (Fig. 6), the collision results in the creation of three pairs of small kinks, and increasing the initial velocity to $v_{0k}=-v_{0ak}=0.9$ at the same $\lambda$ (Fig. 6) leads to the creation of four pairs of small kinks. These results demonstrate that both higher $\lambda$ values and larger initial velocities contribute to the production of more small kinks, as the increased energy available in the collision can overcome the energy barriers required to create additional small kinks.
The creation of small kinks in large kink-antikink collisions can be understood through energy conservation. Large kinks, having a greater mass ($M>m$, where $m$ is the mass of a small kink), possess more kinetic energy for the same initial velocity. During the collision, this kinetic energy is partially converted into the potential energy required to create small kinks, with the number of small kinks produced depending on the available energy, which is influenced by both $\lambda$ and the initial velocity. The absence of vibrational modes in large kinks, as noted earlier, means that the energy exchange during the collision is primarily between translational modes and the potential energy of the system, facilitating the creation of small kinks rather than energy storage in internal modes.
@@ IV Conclusion
Our findings highlight the significant influence of the parameter $\lambda$ on the scattering dynamics of kinks, mediating the transition between different kink scattering regimes and influencing the production of new kinks. The absence of vibrational modes in large kinks simplifies their dynamics, while the presence of such modes in small kinks introduces complex phenomena like resonance and bion formation. These results contribute to a deeper understanding of kink interactions in nonlinear systems and may have implications for applications in fields such as condensed matter physics, where topological defects such as kinks may play a critical role. As we know, magnetic materials at the nanometric scale may produce localized magnetic configurations in the form of kinks or domain walls that are of current interest in applications of practical use, such as the implementation of digital devices MM01 . In the case of kinks that change their internal profile, it is of interest to recall the investigation developed before in MM02 , where the authors studied the possibility that in a nanowire, two distinct types of kinks (known in condensed matter as the Boch wall and the Néel wall) can switch their internal profile from Bloch to Néel wall under modification of the width and thickness of the nanowire. Future studies could explore the effects of additional parameters, such as external fields or temperature, on kink scattering, as well as the potential for experimental verification of these phenomena in physical systems.

### v2

@@ III.2 Large Kinks Scattering
Fig. 6 illustrates the scattering of large kinks for different values of $\lambda$ and initial velocities, with initial positions $X_{0k}=-X_{0ak}=-10$ in all cases. In Fig. 6, for $\lambda=0.2$ and $v_{0k}=-v_{0ak}=0.1$, the collision results in the creation of two small kinks, which move away symmetrically after the interaction. Increasing $\lambda$ to 0.6 with the same initial velocity ($v_{0k}=-v_{0ak}=0.1$), as shown in Fig. 6, also leads to the creation of two small kinks, but with higher velocities, indicating that a larger $\lambda$ enhances the energy transfer to the resulting small kinks. For $\lambda=0.8$ and $v_{0k}=-v_{0ak}=0.1$ (Fig. 6), the collision produces two pairs of small kinks, suggesting that higher $\lambda$ values facilitate the creation of additional small kinks due to changes in the potential $U_{\lambda}(\chi)$. A similar outcome is observed in Fig. 6 for $\lambda=0.6$ but with a higher initial velocity ($v_{0k}=-v_{0ak}=0.9$), where two pairs of small kinks are created, highlighting the role of initial velocity in increasing the number of small kinks produced. For $\lambda=0.9$ and $v_{0k}=-v_{0ak}=0.2$ (Fig. 6), the collision results in the creation of three pairs of small kinks, and increasing the initial velocity to $v_{0k}=-v_{0ak}=0.9$ at the same $\lambda$ (Fig. 6) leads to the creation of four pairs of small kinks. As expected, the results depicted in Figure 6 show that both higher $\lambda$ values and larger initial velocities directly contribute to the production of more small kinks, as the increased energy available in the collision can overcome the energy barriers required to create additional small kinks.
The creation of small kinks in large kink-antikink collisions can be understood through energy conservation. Large kinks, having a greater mass ($M>m$, where $m$ is the mass of a small kink), possess more kinetic energy for the same initial velocity. During the collision, this kinetic energy is partially converted into the potential energy required to create small kinks, with the number of small kinks produced depending on the available energy, which is influenced by both $\lambda$ and the initial velocity. The absence of vibrational modes in large kinks, as noted earlier, means that the energy exchange during the collision is primarily between translational modes and the potential energy of the system, facilitating the creation of small kinks rather than energy storage in internal modes. A quantitative investigation dealing with the number of small kinks created in terms of $\lambda$ and initial velocity $v_{i}$ of the kink and antikink system described in Eq. (31) is also of current interest, since it may be further connected with quantum effects for kink creation, as recently investigated in Ref. vasha . This motivation requires adding other fields, which is out of the scope of the present work and will be left to a future investigation.
@@ IV Conclusion
Our findings highlight the significant influence of the parameter $\lambda$ on the scattering dynamics of kinks, mediating the transition between different kink scattering regimes and influencing the production of new kinks. The absence of vibrational modes in large kinks simplifies their dynamics, while the presence of such modes in small kinks introduces complex phenomena like resonance and bion formation. These results contribute to a deeper understanding of kink interactions in nonlinear systems and may have implications for applications in fields such as condensed matter physics, where topological defects such as kinks may play a critical role. It is well known that magnetic materials at the nanometric scale may produce localized magnetic configurations in the form of kinks or domain walls that are of current interest in applications of practical use, such as the implementation of digital devices MM01 . In the case of kinks that change their internal profile, it is of interest to recall the investigation developed before in MM02 , where the authors studied the possibility that in a nanowire, two distinct types of kinks (known in condensed matter as the Boch wall and the Néel wall) can switch their internal profile from Bloch to Néel wall under modification of the width and thickness of the nanowire. Future studies could explore the effects of additional parameters, such as external fields or temperature, on kink scattering, as well as the potential for experimental verification of these phenomena in physical systems.